# Doubly Optimal Policy Evaluation for Reinforcement Learning

### v1

@@ B.1 GridWorld
For a Gridworld with size $n$, its width, height, and time horizon $T$ are all set to $n$. The number of states in this Gridworld environment scales cubically with $n$, offering a suitable tool to test algorithm scalability. We choose Gridworld with $n^{3}=1,000$ and $n^{3}=27,000$, which are the largest Gridworld environment tested among related works (Jiang and Li,, 2016; Hanna et al.,, 2017; Liu and Zhang,, 2024). There are four possible actions: up, down, left, and right. After taking an action, the agent has a $0.9$ probability of moving accordingly and a $0.1$ probability of moving uniformly at random. If the agent runs into a boundary, the agent stays in its current location. The reward function $r(s,a)$ is randomly generated. We consider $30$ randomly generated target policies. The ground truth policy performance is estimated using the on-policy Monte Carlo method by running each target policy for $10^{6}$ episodes. We test two different sizes of the Gridworld with a number of $1,000$ and $27,000$ states. The offline dataset of both environments contains $1,000$ episodes generated by a set of random policies. To learn functions $q_{\pi,t}$ and $u_{\pi,t}$, we split the offline data into a training set and a test set. We tune all hyperparameters offline based on Fitted Q-learning loss on the test set. We choose a one-hidden-layer neural network and test the neural network size with $[64,128,256]$ and choose $64$ as the final size. We test the learning rate for Adam optimizer with $[1\text{e}^{-5},1\text{e}^{-4},1\text{e}^{-3},1\text{e}^{-2}]$ and choose to use the default learning rate $1\text{e}^{-3}$ as learning rate for Adam optimizer (Kingma and Ba,, 2015). All benchmark algorithms are learned using their reported hyperparameters (Jiang and Li,, 2016; Liu and Zhang,, 2024). Each policy has 30 independent runs, resulting in $30\cdot 30=900$ total runs. Therefore, each curve in Figure 1 is averaged from 900 different runs over a wide range of policies indicating strong statistical significance.
@@ B.2 MuJoCo
MuJoCo is a physics engine containing various stochastic environments, where the goal is to control a robot to achieve different behaviors such as walking, jumping, and balancing. Environments in Figure 3 from the left to the right are Ant, Hopper, InvertedDoublePendulum, InvertedPendulum, and Walker. We construct $30$ policies in each environment (resulting a total of $150$ policies) with a wide range of performance generated by the proximal policy optimization (PPO) algorithm (Schulman et al.,, 2017) with the default PPO implementation in Huang et al., (2022). We set each MuJoCo environment to have a fixed time horizon $100$ in OpenAI Gymnasium (Towers et al.,, 2024). Since our methods are designed for discrete action space, we discretize the first dimension of MuJoCo action space. The remaining dimensions are controlled by the PPO policies and are deemed as part of the environment. The offline dataset of each environment contains $1,000$ episodes generated by a set of policies with various performances. Functions $q_{\pi,t}$ and $u_{\pi,t}$ are learned the same way as in Gridworld environments. Our algorithm is robust on hyperparameters. All hyperparameters in Algorithm 1 are tuned offline and are the same across all MuJoCo and Gridworld experiments. Each policy in MuJoCo also has 30 independent runs, resulting in $30\cdot 30=900$ total runs. Therefore, each curve in Figure 2 and each number in Table 2 are averaged from 900 different runs over a wide range of policies indicating strong statistical significance.

### v2

@@ B.1 GridWorld
For a Gridworld with size $n$, we set its width, height, and time horizon all to be $n$. The number of states in this Gridworld environment scales cubically with $n$, offering a suitable tool to test algorithm scalability. We choose Gridworld with $n^{3}=1,000$ and $n^{3}=27,000$, the largest Gridworld environment tested among related works (Jiang and Li,, 2016; Hanna et al.,, 2017; Liu and Zhang,, 2024). There are four possible actions: left, right, up, and down. After the agent takes an action, it has a probability of $0.9$ to move accordingly and a probability of $0.1$ to move uniformly at random. If runs into a boundary, the agent stays in its current location. The reward function $r(s,a)$ is randomly generated. We consider $30$ randomly generated target policies. We generate the ground truth policy performance using the on-policy Monte Carlo method, running each target policy for $10^{6}$ episodes. We test two different environment sizes of the Gridworld, one with $1,000$ states and $27,000$ states. The offline dataset of both environments contains $1,000$ episodes generated by a set of random policies. To learn functions $q_{\pi,t}$ and $u_{\pi,t}$, we split the offline data into a training set and a test set. We tune all hyperparameters offline based on Fitted Q-learning loss on the test set. We choose a one-hidden-layer neural network and test the neural network size with $[64,128,256]$ and choose $64$ as the final size. We test the learning rate for Adam optimizer with $[1\text{e}^{-5},1\text{e}^{-4},1\text{e}^{-3},1\text{e}^{-2}]$ and choose to use the default learning rate $1\text{e}^{-3}$ as learning rate for Adam optimizer (Kingma and Ba,, 2015). All benchmark algorithms are learned using their reported hyperparameters (Jiang and Li,, 2016; Liu and Zhang,, 2024). Each policy has 30 independent runs, resulting in $30\cdot 30=900$ total runs. Therefore, each curve in Figure 1 is averaged from 900 different runs over a wide range of policies, showing a strong statistical significance.
@@ B.2 MuJoCo
MuJoCo is a physics engine containing various stochastic environments, where the goal is to control a robot to achieve different behaviors such as walking, jumping, and balancing. Environments in Figure 3 from the left to the right are Ant, Hopper, InvertedDoublePendulum, InvertedPendulum, and Walker. We construct $30$ policies in each environment (resulting a total of $150$ policies), incorporating a wide range of performance generated by the proximal policy optimization (PPO) algorithm (Schulman et al.,, 2017). We use the the default PPO implementation in Huang et al., (2022). We set each MuJoCo environment to have a fixed time horizon $100$ in OpenAI Gymnasium (Towers et al.,, 2024). As our methods are designed for discrete action space, we discretize the first dimension of MuJoCo action space. The remaining dimensions are controlled by the PPO policies, and they are deemed as part of the environment. The offline dataset of each environment contains $1,000$ episodes generated by a set of policies with various performances. Functions $q_{\pi,t}$ and $u_{\pi,t}$ are learned the same way as in Gridworld environments. Our algorithm is robust on hyperparameters. All hyperparameters in Algorithm 1 are tuned offline and are the same across all MuJoCo and Gridworld experiments. Each policy in MuJoCo also has 30 independent runs, resulting in $30\cdot 30=900$ total runs. Therefore, each curve in Figure 2 and each number in Table 2 are averaged from 900 different runs over a wide range of policies indicating strong statistical significance.